# Not Just How Much, But Where: Decomposing Epistemic Uncertainty into Per-Class Contributions

### v1

@@ C.1 Monte Carlo Prediction Statistics
with Pearson correlation $\rho_{ij}=\Sigma_{ij}/(\sigma_{i}\,\sigma_{j})$. A strongly negative $\rho_{ij}$ indicates that probability mass flows systematically from class $j$ to class $i$ across forward passes—i.e. the model actively confuses the two classes.
@@ C.2 Scalar Uncertainty Baselines
lies in $0,\log K]$ and captures total predictive uncertainty—both epistemic and aleatoric. It cannot distinguish a model that lacks training data from one facing a genuinely ambiguous input; nor does it differentiate clinically benign confusions (e.g. Grade 0 vs. Grade 1) from dangerous ones (Grade 0 vs. Grade 3).
@@ C.8 Evaluation Metrics for Selective Prediction
$\mathrm{Critical\;FNR}(c)\;=\;\frac{\sum_{n\in\mathrm{kept}(c)}\mathbb{1}[y_{n}\in\mathcal{C}]\cdot\mathbb{1}[\hat{y}_{n}\notin\mathcal{C}]}{\sum_{n\in\mathrm{kept}(c)}\mathbb{1}[y_{n}\in\mathcal{C}]}\,.$ (49)
@@
$\mathrm{Critical\;Err}(c)\;=\;\frac{\sum_{n\in\mathrm{kept}(c)}\mathbb{1}[y_{n}\in\mathcal{C}]\cdot\mathbb{1}[\hat{y}_{n}\neq y_{n}]}{\sum_{n\in\mathrm{kept}(c)}\mathbb{1}[y_{n}\in\mathcal{C}]}\,.$ (50)
@@
$\mathrm{Accuracy}(c)=\frac{\sum_{n\in\mathrm{kept}(c)}\mathbb{1}[\hat{y}_{n}=y_{n}]}{|\mathrm{kept}(c)|},\qquad\mathrm{Macro\;F1}(c)=\frac{1}{K}\sum_{k=0}^{K-1}F_{1,k}\,.$ (52)
@@ D.9 Error Signatures
Table 11: Error epistemic signatures: $\mathbb{E}[C_{k}\mid y{=}i,\hat{y}{=}j]$ for key error types. Errors with nearly identical scalar MI exhibit strikingly different $C_{k}$ profiles, revealing distinct confusion pathways that suggest different remediation strategies.
@@
The key interpretability result is that errors with similar scalar MI have distinctive $C_{k}$ fingerprints. Catastrophic misses (true Grade 3 predicted as Grade 0, n=107n{=}107) and severity underestimates (true Grade 3 predicted as Grade 2, n=220n{=}220) have nearly identical scalar MI ($0.024$ vs. $0.027$ nats) but strikingly different $C_{k}$ signatures. The catastrophic miss concentrates epistemic mass on $C_{2}$: the model does not jump directly from severe to healthy, but routes through moderate DR, identifying Grade 2 as the “bottleneck” of confusion. The severity underestimate instead elevates $C_{0}$, indicating the model’s primary doubt is whether the image might be healthy. These qualitatively different failure modes, invisible to any scalar metric, suggest distinct intervention strategies: additional moderate-vs-severe training data for the first, and better healthy-vs-severe discrimination for the second.
@@ D.11 Full Bootstrap Results: MC Dropout
CBEC reduces AUSC by 53.6% relative to MI ($0.197$ vs. $0.425$) and wins 100% of bootstrap samples. Scalar entropy ($0.249$) outperforms all per-class epistemic methods including MI ($0.425$) under MC dropout; this is because entropy captures total uncertainty (aleatoric + epistemic), and under MC dropout’s coarser posterior, the aleatoric component carries complementary signal that pure epistemic measures miss. $C_{\text{crit\_max}}$ still outperforms its unnormalised counterpart Sale_EU_crit ($0.419$ vs. $0.463$), confirming that the $1/\mu_{k}$ normalisation provides benefit even when the Taylor approximation is degraded.

### v2

@@ C.1 Monte Carlo Prediction Statistics
with Pearson correlation $\rho_{ij}=\Sigma_{ij}/(\sigma_{i}\,\sigma_{j})$. A strongly negative $\rho_{ij}$ indicates that probability mass flows systematically from class $j$ to class $i$ across forward passes, i.e. the model actively confuses the two classes.
@@ C.2 Scalar Uncertainty Baselines
lies in $0,\log K]$ and captures total predictive uncertainty, both epistemic and aleatoric. It cannot distinguish a model that lacks training data from one facing a genuinely ambiguous input; nor does it differentiate clinically benign confusions (e.g. Grade 0 vs. Grade 1) from dangerous ones (Grade 0 vs. Grade 3).
@@ C.8 Evaluation Metrics for Selective Prediction
$\mathrm{Critical\;FNR}(c)\;=\;\frac{\sum_{n\in\mathrm{kept}(c)}\mathbf{1}[y_{n}\in\mathcal{C}]\cdot\mathbf{1}[\hat{y}_{n}\notin\mathcal{C}]}{\sum_{n\in\mathrm{kept}(c)}\mathbf{1}[y_{n}\in\mathcal{C}]}\,.$ (49)
@@
$\mathrm{Critical\;Err}(c)\;=\;\frac{\sum_{n\in\mathrm{kept}(c)}\mathbf{1}[y_{n}\in\mathcal{C}]\cdot\mathbf{1}[\hat{y}_{n}\neq y_{n}]}{\sum_{n\in\mathrm{kept}(c)}\mathbf{1}[y_{n}\in\mathcal{C}]}\,.$ (50)
@@
$\mathrm{Accuracy}(c)=\frac{\sum_{n\in\mathrm{kept}(c)}\mathbf{1}[\hat{y}_{n}=y_{n}]}{|\mathrm{kept}(c)|},\qquad\mathrm{Macro\;F1}(c)=\frac{1}{K}\sum_{k=0}^{K-1}F_{1,k}\,.$ (52)
@@ D.9 Error Signatures
Table 11: Error epistemic signatures: $\mathbb{E}[C_{k}\mid y{=}i,\hat{y}{=}j]$ for key error types. Errors with nearly identical scalar MI exhibit different $C_{k}$ profiles, revealing distinct confusion pathways that suggest different remediation strategies.
@@
The key interpretability result is that errors with similar scalar MI have distinctive $C_{k}$ fingerprints. Catastrophic misses (true Grade 3 predicted as Grade 0, n=107n{=}107) and severity underestimates (true Grade 3 predicted as Grade 2, n=220n{=}220) have nearly identical scalar MI ($0.024$ vs. $0.027$ nats) but quite different $C_{k}$ signatures. The catastrophic miss concentrates epistemic mass on $C_{2}$: the model does not jump directly from severe to healthy, but routes through moderate DR, identifying Grade 2 as the “bottleneck” of confusion. The severity underestimate instead elevates $C_{0}$, indicating the model’s primary doubt is whether the image might be healthy. These qualitatively different failure modes, invisible to any scalar metric, suggest distinct intervention strategies: additional moderate-vs-severe training data for the first, and better healthy-vs-severe discrimination for the second.
@@ D.11 Full Bootstrap Results: MC Dropout
CBEC reduces AUSC by 53.6% relative to MI ($0.197$ vs. $0.425$) and wins 100% of bootstrap samples. Scalar entropy ($0.249$) outperforms all epistemic methods including MI ($0.425$) under MC dropout; this is because entropy captures total uncertainty (aleatoric + epistemic), and under MC dropout’s coarser posterior, the aleatoric component carries complementary signal that pure epistemic measures miss. $C_{\text{crit\_max}}$ still outperforms its unnormalised counterpart Sale_EU_crit ($0.419$ vs. $0.463$), confirming that the $1/\mu_{k}$ normalisation provides benefit even when the Taylor approximation is degraded.